OpenStudy (howard-wolowitz):

What are the conditions for using the normal approximation for a sampling distribution of proportions?

OpenStudy (mathmale):

Hello, Howard! I Googled "normal approximations from sample proportions" and came up with the following result, among many: https://onlinecourses.science.psu.edu/stat200/node/43 In particular, have a look at this: np≥10 and n(1−p)≥10. n=sample size p=sample proportion

OpenStudy (mathmale):

Take n and p from the problem statement and determine whether or not the two inequalities shown above are true for those values.

OpenStudy (zarkon):

One way to do this... if \(p\pm3\sqrt{pq/n}\) lie within (0,1) then \(0<p-3\sqrt{pq/n}\) and \(p+3\sqrt{pq/n}<1\) this is equivalent to \(n>9\left(\dfrac{p}{q}\right)\) and \(n>9\left(\dfrac{q}{p}\right)\) so you want \[n>9\left(\frac{\max(p,q)}{\min(p,q)}\right)\]

OpenStudy (howard-wolowitz):

Ok guys, but i dont understand how i wrtie out these answers in sentence forms

OpenStudy (welshfella):

sorry cant help with this one